What is the electric field strength at a point in space 4 m due east of a particle with a charge of +8 μC? What

is the direction of the field here?



1
Expert's answer
2021-04-15T10:31:31-0400

Answer

Electric field at a point in space can be given as

"E=\frac{kq}{r^2}"


So

"E=\frac{(9\times10^9) (8\times10^{-6}) }{4^2}\\E=4500N/C"


Direction will be towards east.